# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v1

@@ 4.1 Optimization Strategy
1:2D pose $\mathbf{J}_{j}^{2D}$, pre-trained model $f^{*}$ with fixed parameters, projection matrix $\mathbf{P}$, optimization steps $T$
2:3D pose $\hat{\mathbf{J}}_{j}^{3D}$
3:Initialization: $\mathbf{z}_{j}=\mathbf{J}_{j}^{2D}$, $iter=0$
4:$\hat{\boldsymbol{\mu}}_{j}=f_{\boldsymbol{\mu}}^{*}(\mathbf{J}_{j}^{2D}),\hat{\boldsymbol{s}}_{j}=f_{\boldsymbol{s}}^{*}(\mathbf{J}_{j}^{2D})$
5:while $iter<T$ do
6:     $\hat{\mathbf{J}}_{j}^{3D}=f_{\boldsymbol{\mu}}^{*}(\mathbf{z}_{j})$
7:     $L_{P}^{opt}=\|\hat{\mathbf{J}}_{j}^{3D}\cdot\mathbf{P}-\mathbf{J}_{j}^{2D}\|_{2}$
8:     $L_{U}^{opt}=\left\|\hat{\boldsymbol{\mu}}_{j}-\hat{\mathbf{J}}_{j}^{3D}\right\|^{2}/(2\cdot\exp(\hat{\boldsymbol{s}}_{j}))$
9:     $L_{total}^{opt}=\lambda_{P}\cdot L_{P}^{opt}+\lambda_{U}\cdot L_{U}^{opt}$
10:     $\mathbf{z}_{j}\leftarrow Adam(\mathbf{z}_{j},\nabla L_{total}^{opt})$
11:     $iter=iter+1$
12:end while
13:$\hat{\mathbf{J}}_{j}^{3D}=f_{\boldsymbol{\mu}}^{*}(\mathbf{z}_{j})$
14:return $\hat{\mathbf{J}}_{j}^{3D}$
[…]
@@
SimpleBaseline (ICCV’2017) [21] 51.8 56.2 58.1 59.0 69.5 78.4 55.2 58.1 74.0 94.6 62.3 59.1 65.1 49.5 52.4 62.9
VideoPose3D (CVPR’2019) [23] 47.1 50.6 49.0 51.8 53.6 61.4 49.4 47.4 59.3 67.4 52.4 49.5 55.3 39.5 42.7 51.8
LCN (ICCV’2019) [4] 46.8 52.3 44.7 50.4 52.9 68.9 49.6 46.4 60.2 78.9 51.2 50.0 54.8 40.4 43.3 52.7
[…]
SRNet (ECCV’2020) [37] 44.5 48.2 47.1 47.8 51.2 56.8 50.1 45.6 59.9 66.4 52.1 45.3 54.2 39.1 40.3 49.9
GraphSH (CVPR’2021) [35] 45.2 49.9 47.5 50.9 54.9 66.1 48.5 46.3 59.7 71.5 51.4 48.6 53.9 39.9 44.1 51.9
[…]
GraFormer (CVPR’2022) [45] 45.2 50.8 48.0 50.0 54.9 65.0 48.2 47.1 60.2 70.0 51.6 48.7 54.1 39.7 43.1 51.8
UGRN (AAAI’2023) [16] 47.9 50.0 47.1 51.3 51.2 59.5 48.7 46.9 56.0 61.9 51.1 48.9 54.3 40.0 42.9 50.5
MLP-JCG (TMM’2023) [30] 43.8 46.7 46.9 48.9 50.3 60.1 45.7 43.9 56.0 73.7 48.9 48.2 50.9 39.9 41.5 49.7
GUMLP (Ours) 44.3 49.8 44.7 49.3 51.4 58.1 47.3 45.6 58.2 63.7 51.0 47.6 53.4 38.1 40.4 49.5
GUMLP + UAO (Ours) 42.3 47.7 42.8 47.8 48.7 55.8 45.8 42.4 55.2 60.9 48.6 45.9 50.2 36.0 38.2 47.2
@@
SimpleBaseline (ICCV’2017) [21] 37.7 44.4 40.3 42.1 48.2 54.9 44.4 42.1 54.6 58.0 45.1 46.4 47.6 36.4 40.4 45.5
STGCN (ICCV’2019) [2] 33.4 39.0 33.8 37.0 38.1 47.3 39.5 37.3 43.2 46.2 37.7 38.0 38.6 30.4 32.1 38.1
LCN (ICCV’2019) [4] 36.3 38.8 29.7 37.8 34.6 42.5 39.8 32.5 36.2 39.5 34.4 38.4 38.2 31.3 34.2 36.3
MGCN (ICCV’2021) [47] - - - - - - - - - - - - - - - 37.4
GraFormer (CVPR’2022) [45] 32.0 38.0 30.4 34.4 34.7 43.3 35.2 31.4 38.0 46.2 34.2 35.7 36.1 27.4 30.6 35.2
MLP-JCG (TMM’2023) [30] 29.1 36.0 30.4 33.8 35.5 46.5 35.3 31.2 39.2 48.8 33.9 35.2 35.8 26.9 29.4 35.1
@@
4.2 Projection Constraint
Projection constraint is commonly used in optimization-based methods, such as SPIN [15] and EFT [12]. For a given 3D pose $\mathbf{J}^{3D}$, its 2D projection $\mathbf{J}^{2D}$ should conform to the principles of perspective projection [33], which can be expressed as follow:
$\mathbf{J}^{3D}\cdot\mathbf{P}=\mathbf{J}^{2D},$ (4)
where the projection matrix $\mathbf{P}\in\mathbb{R}^{3\times 2}$ contains the intrinsic parameters of the camera. In real scenarios, the intrinsic parameters of the camera can be obtained from camera specifications or can be inferred solely from the input images [34]. During the testing-time optimization, the 2D projection $\hat{\mathbf{J}}^{2D}=\hat{\mathbf{J}}^{3D}\cdot\mathbf{P}$ of the predicted 3D pose $\hat{\mathbf{J}}^{3D}$ is expected to closely align with the given 2D pose $\mathbf{J}^{2D}$. A large error between them indicates that the estimated 3D pose $\hat{\mathbf{J}}^{3D}$ may be problematic and needs to be corrected. The formulation of projection constraint can be expressed as:
$L^{opt}_{P}=||\hat{\mathbf{J}}^{2D}-\mathbf{J}^{2D}||_{2}.$ (5)
The purpose of the projection constraint is to enforce the relationship between the estimated 3D pose $\hat{\mathbf{J}}^{3D}$ and its corresponding 2D projection $\hat{\mathbf{J}}^{2D}$ to follow the geometric principles of perspective projection, thereby improving the accuracy of the pose estimation.
4.3 Uncertainty Constraint
@@
During the testing process, for the $j$-th 2D input, the pre-trained model allows us to acquire the initial 3D result $\hat{\boldsymbol{\mu}}_{k}$ and uncertainty $\hat{\boldsymbol{s}}_{k}$ associated with the $k$-th estimated 3D joint. We can explicitly express the uncertainty constraint as:
$\mathcal{L}_{U,j}^{opt}=\frac{1}{2K}\sum_{k=1}^{K}\left\|\hat{\boldsymbol{\mu}}_{j,k}-f_{\boldsymbol{\mu}}^{*}(\mathbf{z})_{j,k}\right\|^{2}\exp(-\hat{\boldsymbol{s}}_{j,k}),$ (6)
where $f^{*}$ denotes the pre-traind model, $f_{\boldsymbol{\mu}}$ denotes the predicted 3D result. This constraint enables joints with higher uncertainty less constrained by the results of the initial estimation, allowing more room for optimization. On the other hand, the joint with lower uncertainty means that the network is confident about the result, so limit it from further drifting.
5 Experiments
5.1 Datasets and Evaluation Metrics
@@
Human3.6M is the most representative benchmark for 3D human pose estimation. It contains 3.6 million video frames captured from four synchronized cameras at 50Hz in an indoor environment. There are 11 professional actors performing 17 actions, such as greeting, phoning, and sitting. Following previous works [23, 2], we train our model on five subjects (S1, S5, S6, S7, S8) and test it on two subjects (S9 and S11). The performance is evaluated by two common metrics: MPJPE (Mean Per-Joint Position Error), termed Protocol 1, is the mean Euclidean distance between the predicted joints and the ground truth in millimeters. P-MPJPE, termed Protocol 2, is the MPJPE after rigid alignment with the ground truth in translation, rotation, and scale.
@@
MPI-INF-3DHP is a more challenging 3D pose dataset that contains both indoor and complex outdoor scenes. There are 8 actors performing 8 actions from 14 camera views, which cover a greater diversity of poses. Its test set consists of three different scenes: studio with green screen (GS), studio without green screen (noGS), and outdoor scene (Outdoor). Following [47, 38, 23], we use Percentage of Correct Keypoints (PCK) with a threshold of 150mm and the Area Under Curve (AUC) for a range of PCK thresholds for evaluation.
5.2 Implementation Details
@@
We implement our approach with PyTorch, training and testing it on one NVIDIA RTX 3090 GPU. Following [47, 2], we use 2D pose detected by CPN [3] for Human3.6M, and ground truth 2D pose for MPI-INF-3DHP. Horizontal flipping is used as data augmentation following [47, 2, 37]. During the training phase, we train a deep neural network for deep prior. The proposed GUMLP is designed by stacking GCN layers and UMLP modules for $N=4$ times. Adam optimizer is adopted with an initial learning rate of 0.001 and a decay factor of 0.95 per epoch. For the testing time optimization, the hyperparameters for projection loss and uncertainty loss are chosen as $\lambda_{P}=1$ and $\lambda_{U}=0.005$, respectively.
Table 2: Quantitative comparisons with state-of-the-art methods on the MPI-INF-3DHP dataset.
@@
GUMLP (Ours) 87.6 86.1 87.1 87.0 54.3
GUMLP + UAO (Ours) 89.2 87.7 87.9 88.2 54.5
5.3 Comparison with State-Of-the-Art Methods
Human3.6M. The comparison results with other state-of-the-art methods under Protocol 1 are shown in Table 1. Following [47, 2, 35], the 2D pose detected by CPN [3] is used as input for training and testing. It is worth mentioning that the previous state-of-the-art method, MGCN [47], incorporated a refinement module to enhance performance. However, even without this module, our baseline model GUMLP achieves a comparable accuracy, with only a marginal difference of 0.1mm. Furthermore, by integrating the Uncertainty-Aware testing-time Optimization (UAO) framework, which leverages both projection and uncertainty constraints, the performance of GUMLP is significantly improved (from 49.5mm to 47.2mm, a relative 4.6% improvement). Notably, GUMLP with UAO framework outperforms MGCN [47] by 4.5% (49.4mm vs. 47.2mm). For Protocol 2, we also obtain the best overall results (39.0mm, see supplemental material). These results validate the effectiveness of the designed 2D-to-3D lifting network (GUMLP) and the UAO framework. To further explore the lower bounds of our approach, we compare our approach with previous state-of-the-art methods with ground truth 2D pose as input. As shown in Table 1 (bottom), we achieve state-of-the-art performance, outperforming all other methods. Furthermore, we observe that the performance of GUMLP can still be boosted with the UAO framework, even when the latent state $\mathbf{z}$ is initialized with the ground truth 2D pose. This is due to the insufficient capacity of the pre-trained GUMLP, and the use of projection constraint and uncertainty constraint during the optimization process can effectively enhance the quality of generated 3D poses.
@@ 5.3 Comparison with State-Of-the-Art Methods
MPI-INF-3DHP. To evaluate the generalization ability of the proposed UAO framework, we further compare our approach against previous state-of-the-art methods on cross-dataset scenarios. As shown in Table 2, GUMLP obtains the best results in all scenes and all metrics, consistently surpassing other methods. After utilizing the UAO framework, the performance can be further improved in both indoor and outdoor scenarios. This suggests that the designed testing-time optimization process can effectively enhance the model’s ability to generalize to unknown actions and datasets.
Table 3: Ablation study on proposed constraints in 5 iterations step using BOA framework.
Pre-trained Projection Uncertainty MPJPE $\downarrow$
Model Constraint Constraint
49.5
48.2
47.2
5.4 Ablation Study

### v2

@@ 9 $iter=iter+1$
TABLE I: Quantitative comparison with the state-of-the-art methods on Human3.6M under Protocol 1. Top-table: 2D pose detected by cascaded pyramid network (CPN) [29] is used as input. Bottom-table: Ground truth 2D pose is used as input. The top two best results for each action are highlighted in bold and underlined, respectively.
@@
SimpleBaseline [4] 51.8 56.2 58.1 59.0 69.5 78.4 55.2 58.1 74.0 94.6 62.3 59.1 65.1 49.5 52.4 62.9
VideoPose3D [26] 47.1 50.6 49.0 51.8 53.6 61.4 49.4 47.4 59.3 67.4 52.4 49.5 55.3 39.5 42.7 51.8
LCN [58] 46.8 52.3 44.7 50.4 52.9 68.9 49.6 46.4 60.2 78.9 51.2 50.0 54.8 40.4 43.3 52.7
SRNet [59] 44.5 48.2 47.1 47.8 51.2 56.8 50.1 45.6 59.9 66.4 52.1 45.3 54.2 39.1 40.3 49.9
GraphSH [32] 45.2 49.9 47.5 50.9 54.9 66.1 48.5 46.3 59.7 71.5 51.4 48.6 53.9 39.9 44.1 51.9
MGCN [7] 45.4 49.2 45.7 49.4 50.4 58.2 47.9 46.0 57.5 63.0 49.7 46.6 52.2 38.9 40.8 49.4
GraFormer [60] 45.2 50.8 48.0 50.0 54.9 65.0 48.2 47.1 60.2 70.0 51.6 48.7 54.1 39.7 43.1 51.8
UGRN [61] 47.9 50.0 47.1 51.3 51.2 59.5 48.7 46.9 56.0 61.9 51.1 48.9 54.3 40.0 42.9 50.5
MLP-JCG [62] 43.8 46.7 46.9 48.9 50.3 60.1 45.7 43.9 56.0 73.7 48.9 48.2 50.9 39.9 41.5 49.7
DiffPose [63] - - - - - - - - - - - - - - - 49.7
GUMLP (Ours) 44.3 49.8 44.7 49.3 51.4 58.1 47.3 45.6 58.2 63.7 51.0 47.6 53.4 38.1 40.4 49.4
GUMLP + UAO (Ours) 42.3 47.5 42.8 47.8 48.7 55.8 45.8 42.4 55.2 60.9 48.6 45.9 50.2 36.0 38.2 46.7
@@
SimpleBaseline [4] 37.7 44.4 40.3 42.1 48.2 54.9 44.4 42.1 54.6 58.0 45.1 46.4 47.6 36.4 40.4 45.5
LCN [58] 36.3 38.8 29.7 37.8 34.6 42.5 39.8 32.5 36.2 39.5 34.4 38.4 38.2 31.3 34.2 36.3
MGCN [7] - - - - - - - - - - - - - - - 37.4
GraFormer [60] 32.0 38.0 30.4 34.4 34.7 43.3 35.2 31.4 38.0 46.2 34.2 35.7 36.1 27.4 30.6 35.2
MLP-JCG [62] 29.1 36.0 30.4 33.8 35.5 46.5 35.3 31.2 39.2 48.8 33.9 35.2 35.8 26.9 29.4 35.1
DAF-DG [64] - - - - - - - - - - - - - - - 44.4
@@
V-C Uncertainty Constraint
@@
During the testing process, for the $j$-th 2D input, the pre-trained model provides the initial 3D result $\hat{\boldsymbol{\mu}}_{j,k}$ and uncertainty $\hat{\boldsymbol{s}}_{j,k}$ associated with the $k$-th estimated 3D joint. We can explicitly formulate the uncertainty constraint $L_{U,j}^{opt}$ as:
$L_{U,j}^{opt}=\frac{1}{2K}\sum_{k=1}^{K}\left\|\hat{\boldsymbol{\mu}}_{j,k}-f_{\boldsymbol{\mu}}^{*}(\mathbf{z})_{j,k}\right\|^{2}\exp(-\hat{\boldsymbol{s}}_{j,k}),$ (11)
where $f^{*}$ denotes the pre-traind model, $f_{\boldsymbol{\mu}}^{*}$ denotes the predicted 3D result. This constraint enables joints with higher uncertainty less constrained by the results of the initial estimation, allowing more room for optimization. On the other hand, the joint with lower uncertainty means that the network is confident about the result, so limits it from further drifting.
V-D Combined Optimization Loss Function
The overall optimization loss $L_{total,j}^{opt}$ for the $j$-th case is defined as the weighted sum of the projection loss and the uncertainty loss:
Ł$\L _{total,j}^{opt}=\lambda_{P}\,L_{P,j}^{opt}+\lambda_{U}\,L_{U,j}^{opt},$ (12)
where $\lambda_{P}$ and $\lambda_{U}$ are hyperparameters for each loss term.
VI Experiments
VI-A Datasets and Evaluation Metrics
@@
Human3.6M [65] is the most representative benchmark for the estimation of 3D human poses. It contains 3.6 million video frames captured from four synchronized cameras at 50Hz in an indoor environment. There are 11 professional actors performing 17 actions, such as greeting, phoning, and sitting. Following previous works [26, 8], we train our model on five subjects (S1, S5, S6, S7, S8) and test it on two subjects (S9 and S11). The performance is evaluated by two common metrics: MPJPE (Mean Per-Joint Position Error), termed Protocol 1, is the mean Euclidean distance between the predicted joints and the ground truth in millimeters. PA-MPJPE, termed Protocol 2, measures the MPJPE after Procrustes Analysis (PA) [66].
@@
MPI-INF-3DHP [67] is a more challenging 3D pose dataset that contains both indoor and complex outdoor scenes. There are 8 actors performing 8 actions from 14 camera views, which cover a greater diversity of poses. Its test set consists of three different scenes: studio with green screen (GS), studio without green screen (noGS), and outdoor scene (Outdoor). Following [7, 68, 26], we use Percentage of Correct Keypoints (PCK) with a threshold of 150mm and the Area Under Curve (AUC) for a range of PCK thresholds for evaluation.
3DPW [69] is a complex in-the-wild dataset that provides 3D pose and mesh annotations, obtained using hand-held cameras and IMUs. It comprises 60 videos recorded at 30 frames per second. To ensure a fair comparison, we train our model on Human3.6M and test it on the 3DPW training set.
VI-B Implementation Details
@@
We implement our approach with PyTorch, training and testing it on one NVIDIA RTX 3090 GPU. Following [7, 8], we use 2D pose detected by CPN [29] for Human3.6M, and ground truth 2D pose for MPI-INF-3DHP. During the training phase, we train a deep neural network for deep prior. The proposed GUMLP is designed by stacking GCN layers and UMLP modules for $N=3$ times. For the testing time optimization, the hyperparameters for projection loss and uncertainty loss are chosen as $\lambda_{P}=1$ and $\lambda_{U}=0.005$, respectively.
TABLE II: Quantitative comparison with the state-of-the-art methods on Human3.6M under Protocol 2. SMPLify-X [40] uses 2D pose detected by OpenPose [70] as input, whereas other methods use 2D poses detected by cascaded pyramid network (CPN) [29]. $\S$ denotes the methods that use refinement module [7, 8]. The top two best results for each action are highlighted in bold and underlined, respectively.
SMPLify-X [40] - - - - - - - - - - - - - - - 75.9
SimpleBaseline [4] 39.5 43.2 46.4 47.0 51.0 56.0 41.4 40.6 56.5 69.4 49.2 45.0 49.5 38.0 43.1 47.7
VideoPose3D [26] 36.0 38.7 38.0 41.7 40.1 45.9 37.1 35.4 46.8 53.4 41.4 36.9 43.1 30.3 34.8 40.0
LCN [58] 36.9 41.6 38.0 41.0 41.9 51.1 38.2 37.6 49.1 62.1 43.1 39.9 43.5 32.2 37.0 42.2
STGCN [8]$\S$ 36.8 38.7 38.2 41.7 40.7 46.8 37.9 35.6 47.6 51.7 41.3 36.8 42.7 31.0 34.7 40.2
SRNet [59] 35.8 39.2 36.6 36.9 39.8 45.1 38.4 36.9 47.7 54.4 38.6 36.3 39.4 30.3 35.4 39.4
MGCN [7]$\S$ 35.7 38.6 36.3 40.5 39.2 44.5 37.0 35.4 46.4 51.2 40.5 35.6 41.7 30.7 33.9 39.1
MLP-JCG [62] 33.7 37.4 37.3 39.6 39.8 47.1 33.7 33.8 45.7 60.5 39.7 37.7 40.1 30.1 33.8 39.3
GKONet [71] 35.4 38.8 35.9 40.4 39.6 44.0 36.7 35.4 46.8 53.7 40.9 36.6 42.0 30.6 33.9 39.4
ZEDO [72] - - - - - - - - - - - - - - - 42.1
VI-C Comparison with State-of-the-art Methods
Human3.6M. The comparison results with other state-of-the-art methods under Protocol 1 are shown in Table I. Following [7, 8, 32], the 2D pose detected by CPN [29] is used as input for training and testing. We have observed that our baseline model, GUMLP, produces competitive results. It is worth noting that none of the compared methods employed a post-optimization approach. To further enhance the accuracy of our model predictions, we integrate the Uncertainty-Aware Optimization (UAO) framework, which leverages both projection and uncertainty constraints. This integration significantly improved GUMLP’s performance, reducing the error from 49.4mm to 46.7mm in MPJPE, a relative improvement of 5.5%. For Protocol 2, we also obtain the best overall results, as shown in Table II. These results validate the effectiveness of the designed 2D-to-3D lifting network (GUMLP) and the UAO framework. To further explore the lower bounds of our approach, we compare it with previous state-of-the-art methods with ground truth 2D pose as input. As shown in Table I (bottom), we achieve state-of-the-art performance, outperforming all other methods. Additionally, we observe that the performance of GUMLP can be further enhanced with the UAO framework, even when the latent state $\mathbf{z}$ is initialized with the ground truth 2D pose. This improvement is attributed to the limited capacity of the pre-trained GUMLP and the effective enhancement of 3D pose quality through the application of projection and uncertainty constraints during the optimization process. This indicates hat optimizing the latent state $\mathbf{z}$ aims to refine the 3D pose, rather than simply making $\mathbf{z}$ closer to the ground truth 2D pose.
TABLE III: Quantitative comparisons with state-of-the-art methods on the MPI-INF-3DHP dataset.
@@ VI-C Comparison with State-of-the-art Methods
GUMLP 87.6 85.9 89.3 87.4 54.8
GUMLP + UAO 88.0 86.9 89.8 88.0 55.1
@@
MPI-INF-3DHP. To evaluate the generalization ability of the proposed UAO framework, we further compare our approach against previous state-of-the-art methods on cross-dataset scenarios. As shown in Table III, GUMLP obtains the best results in all scenes and all metrics, consistently surpassing other methods. After utilizing the UAO framework, the performance can be further improved in both indoor and outdoor scenarios. This suggests that the designed testing-time optimization process can effectively enhance the model’s ability to generalize to unknown actions and datasets.
TABLE IV: Cross-domain evaluation on 3DPW. Source dataset: Human3.6M, target dataset: 3DPW.
Method PA-MPJPE $\downarrow$ MPJPE $\downarrow$
ISO [20] 70.8 -
Wang et al. [73] 68.3 109.5
PoseAug [9] 58.5 94.1
PoseDA [74] 55.3 87.7
PMCE [75] 52.3 81.6
AdaptPose [10] 46.5 81.2
ZEDO [72] 42.6 80.9
BOA [17] 49.5 77.2
DynaBOA [16] 40.4 65.5
DAPA [76] 46.5 75.0
CycleAdapt [22] 39.9 64.7
ESCAPE [21] 47.9 79.0
GUMLP (Ours) 42.0 67.5
GUMLP + UAO (Ours) 41.1 60.8
3DPW. We further evaluated our model’s performance on the 3DPW dataset, which features complex outdoor scenes. As shown in Table IV, our baseline model, GUMLP, outperforms previous methods in both PA-MPJPE and MPJPE metrics. Furthermore, the integration of the UAO framework improves performance by reducing MPJPE by 6.7mm. This result highlights the effectiveness of our proposed optimization strategy in outdoor environments.
VI-D Ablation Study
Impact of Model Design. The GUMLP model primarily comprises two components: a GCN layer and a parallel U-shaped multi-layer perception (UMLP) module, as illustrated in Figure 2. To investigate the design choices of GUMLP, experiments were conducted on Human3.6M using the 2D poses extracted by CPN [29] as input. The results are reported in Table V. Our findings indicate that using only the spatial UMLP, which captures multi-scale features, or only the channel UMLP, which captures multi-level features, does not yield the best performance. The optimal results are achieved when both modules are used together, demonstrating that a combination of multi-level and multi-scale features is necessary for optimal performance.
Impact of Model Configurations. We conducted ablation studies by exploring the performance of GUMLP with different hyperparameters. The results are presented in Table VI. We observe that with a fixed number of layers $L$, the model consistently achieves optimal performance with a hidden size $C$ of 512. Doubling or halving the hidden size results in performance degradation.This is because too few parameters lead to underfitting, while too many cause overfitting and hinder convergence. With a hidden size $C$ of 512, we determined that the optimal number of layers is $L=3$ for the best performance. In Table VII, we further explore the performance of Spatial UMLP with different spatial dimension $S_{mid}$ after the up-projection layer. The results show that when $S_{mid}=17*6$, our model achieves the best performance.
TABLE V: Ablation study on different designs of our model.
GCN Channel UMLP Spatial UMLP MPJPE $\downarrow$
50.5
56.2
49.4
TABLE VI: Ablation study on various configurations of GUMLP. $L$ denotes the number of layers, $C$ denotes the hidden size.
L C Params (M) FLOPs (M) MPJPE $\downarrow$
2 256 1.55 25.27 52.1
2 512 5.91 76.44 50.8
2 1024 23.33 256.49 54.0
3 256 1.92 37.25 50.8
3 512 7.32 112.1 49.4
3 1024 28.87 374.61 50.6
4 256 2.29 49.23 50.5
4 512 8.72 147.76 49.6
4 1024 34.41 492.72 51.4
TABLE VII: Ablation study on Spatial UMLP. $L$ denotes the number of layers, $C$ denotes the hidden size. $S_{mid}$ denotes the spatial dimensions after the up-projection layer.
L C $S_{mid}$ Params (M) FLOPs (M) MPJPE $\downarrow$
3 512 17 $\times$ 1 7.21 78.8 50.0
3 512 17 $\times$ 3 7.23 86.7 50.1
3 512 17 $\times$ 6 7.32 112.1 49.4
3 512 17 $\times$ 9 7.45 153.1 50.0
Impact of the Initialization Prior. In Table VIII, we use a template 2D pose to initialize the latent state $\mathbf{z}$, treating this as a condition lacking an effective initialization prior. Rows 1 and 4 correspond to general data-driven approaches, where the 2D pose is directly fed into a fixed network to generate an estimated 3D pose without any further optimization. Comparing rows 1–3 with rows 4–6, it is evident that a good initialization is crucial for achieving favorable convergence. Even the worst-case scenario of using the input 2D pose for initialization performs better than cases without an initialization prior. These results emphasize the significance of an effective initialization strategy in improving overall performance.
Impact of Constraints during Optimization. Setting appropriate targets for test-time optimization is essential for achieving high-quality 3D poses. To this end, we evaluate the effectiveness of two proposed constraints in the optimization process. Table VIII presents the ablation results using the UAO framework on the Human3.6M test set with different constraints. Comparing rows 2–3 with rows 5–6, the results show that the projection constraint significantly enhances the performance of the purely data-driven method. Furthermore, the introduction of the uncertainty constraint further improves the model’s generalization ability, leading to superior performance. These results highlight the effectiveness of both constraints in the optimization process. To further explore the optimization mechanism, we illustrate the performance with different iterations on the test set of Human3.6M, as shown in Figure 4. We observe that using only the projection constraint leads to a rapid decrease in MPJPE, followed by a gradual increase. In contrast, the error can be stabilized at a lower level with the further introduction of the uncertainty constraint, even with more iterations. The main reason is that the projection constraint only enforces consistency in the 2D plane, which is easy to overfit, yielding unrealistic poses in the 3D space. To complement this, the uncertainty constraint helps maintain the output within a plausible 3D manifold. It restricts joints with low uncertainty from deviating too far from the ground truth while allowing joints with high uncertainty to have a larger divergence radius. Experimental results demonstrate the robust adaptability of our method, especially in real-world scenarios where the optimal number of iterations is uncertain.
TABLE VIII: Ablation study on initialization prior and constraints.
Row Initialization Initialization Projection Uncertainty MPJPE $\downarrow$
Index Manner Prior Constraint Constraint
1 2D Template Pose 351.8
2 93.1
3 69.5
4 Input 2D Pose 49.4
5 47.1
6 46.7